# VerifyBench: Benchmarking Reference-based Reward Systems for Large Language Models

### v2

@@ Abstract
VerifyBench: Benchmarking Reference-based Reward Systems
for Large Language Models
Yuchen Yan1,2,thanks: ContributionduringinternshipatMeituanGroup., Jin Jiang2,3, Zhenbang Ren1,4, Yijun Li1, Xudong Cai1,5, Yang Liu2, Xin Xu6, Mengdi Zhang2, Jian Shao1,thanks: Correspondingauthor., Yongliang Shen1,†, Jun Xiao1, Yueting Zhuang1 1Zhejiang University  2Meituan Group  3Peking University 4University of Electronic Science and Technology of China 5Beijing University of Posts and Telecommunications 6The Hong Kong University of Science and Technology {yanyuchen, syl, jshao}@zju.edu.cn
@@ 1 Introduction
In recent years, large language models (LLMs) have exhibited remarkable capabilities, significantly assisting humans across diverse practical domains (DeepSeek-AI et al., 2025b; Grattafiori et al., 2024; Yang et al., 2025). Reinforcement learning from human feedback (RLHF) has been crucial to this progress, with reward models playing a central role by evaluating and scoring model-generated responses to guide training. This approach has led to the development of numerous benchmarks  (Lambert et al., 2024; Liu et al., 2024; Zhou et al., 2025) for systematic reward model evaluation, focusing primarily on pairwise preference judgments between competing responses.
Figure 1: The core distinction between VerifyBench and existing reward benchmarks (Lambert et al., 2024; Liu et al., 2024) is illustrated as follows. Upper panel: Existing reward benchmarks assess the accuracy of a reward system by comparing the ranking of two completions for the same question. Lower panel: In contrast, our proposed VerifyBench evaluates the accuracy of a reward system by determining the correctness of a single completion using a reference answer.
The emergence of specialized large reasoning models (LRMs) (DeepSeek-AI et al., 2025a; Qwen Team, 2024; Kimi Team et al., 2025) such as OpenAI’s o1(OpenAI, 2024) and DeepSeek-R1(DeepSeek-AI et al., 2025a) has fundamentally changed this landscape. These models achieve unprecedented performance on reasoning tasks through specialized reinforcement learning techniques that differ from standard RLHF approaches. A key distinction in training methodologies for LRMs is their reliance on reference-based reward systems, where rewards are assigned based on alignment between model-generated responses and authoritative reference answers. This approach has been implemented variously across leading models, with Deepseek-R1 (DeepSeek-AI et al., 2025a) employs a rule-based reward to prevent reward hacking, whereas models like Seed1.5-Thinking (Seed et al., 2025) adopt model-based reward systems to generate more precise and robust signals.
Despite the widespread adoption of reference-based reward systems in training state-of-the-art reasoning models, a significant gap exists in our ability to evaluate these systems systematically. Current benchmarks focus almost exclusively on preference-based evaluation, assessing rewards on their ability to rank competing responses correctly. This approach fails to capture the requirements of reference-based verification, where responses must be judged against objective ground truths rather than relative preferences. The absence of dedicated benchmarks for reference-based reward systems has limited researchers’ ability to assess, compare, and improve their verification methodologies effectively, potentially impeding progress in reasoning model development.
@@ 2 Preliminaries
$\text{Acc}=\frac{1}{|D|}\sum_{(q,r_{w},r_{l})\in D}\mathbb{I}\left[R_{\varphi}(q,r_{w})>R_{\varphi}(q,r_{l})\right]$ (2)
@@
where $q$ denotes the user-issued query, $gt$ denotes the ground-truth reference answer, $r$ represents the model-generated response, and $\varphi$ encapsulates either the learned parameters of the reward model or the established evaluation criteria used to assess the alignment among $q$, $gt$, and $r$.
@@
$\text{Acc}=\frac{1}{|D|}\sum_{(q,gt,r,y)\in D}\mathbb{I}\left[\mathbb{E}(R_{\varphi}(q,gt,r))=y\right]$ (4)
@@ 3 Benchmark Construction
In this paper, we introduce two benchmarks, VerifyBench and VerifyBench-Hard, to evaluate reference-based reward systems. The VerifyBench benchmark is designed to reflect naturally distributed data, whereas VerifyBench-Hard comprises samples exhibiting high levels of disagreement among models, thereby assessing a model’s ability to provide reliable judgments in ambiguous or challenging scenarios.
@@ 3.1 Construction of VerifyBench
To emulate realistic reinforcement learning (RL) scenarios involving reference-based reward systems, we curate a comprehensive collection of open-source reasoning problems paired with corresponding reference answers. These problems encompass three primary categories, general reasoning, logical reasoning and mathematical reasoning, and are aggregated from 41 distinct sources. A complete list of these data sources is provided in Appendix B.
@@
To comprehensively evaluate model performance across diverse answer formats, we define four canonical answer types: numerical values, algebraic expressions, multiple-choice selections, and free-form strings. Utilizing a general-purpose LLM Llama-3.3-70B-Instruct (Grattafiori et al., 2024), we performed automatic answer-type classification with a prompt(Appendix  C.1). Questions that fall outside these categories, such as proof-based or open-ended prompts, were excluded from further analysis. Following classification, we randomly sampled 2,000 instances per answer type, resulting in a final candidate pool of 8,000 questions.
@@
We subsequently employed 22 widely used open-source and proprietary models (see Appendix E) to generate single-shot completions for the curated set of 8,000 questions, resulting in a total of 176,000 completions. To assign initial correctness labels, we utilized Llama-3.3-70B-Instruct (Grattafiori et al., 2024) within a prompt-based judgment framework. For each question, we randomly selected four completions, two labeled as correct and two labeled as incorrect by the model, and retained them for subsequent human annotation.
@@
We conducted human annotation for the aforementioned questions and their associated completions. The annotation procedure comprised two primary tasks: (1) identifying the most appropriate answer type for each question based on its formulation and corresponding ground-truth answer, and (2) evaluating the correctness of each of the four completions. Each question was independently annotated by at least two annotators. If their annotations were consistent, the labeling was finalized; otherwise, a third annotator resolved disagreements to ensure consistency and finalize the labels.
@@
Following human annotation, we identified notable biases in the models’ predictions regarding both answer types and completion correctness, leading to imbalanced data distributions. To mitigate this issue, we performed controlled downsampling to ensure uniform category-level representation and balanced correctness labels. Specifically, we retained 250 questions per answer type, resulting in a total of 1,000 questions. Each question is paired with exactly two completions, one correct and one incorrect. The resulting dataset, VerifyBench, thus comprises 2,000 well-balanced question-answer-completion-correctness tuples. Detailed statistics for VerifyBench are provided in Table 1.
@@
Statistics of Answer Type
@@
Table 1: Benchmark statistics of VerifyBench and VerifyBench-Hard.
@@
OpenAI/gpt-4o-2024-11-20 94.80 90.20 96.80 90.80 93.15 71.43 65.91 75.35 71.30 72.60
OpenAI/gpt-4o-mini 95.80 89.80 95.80 90.00 92.85 69.05 72.73 74.19 72.17 72.30
meta-llama/Llama-4-Scout-17B-16E-Instruct 94.20 86.80 89.80 89.25 90.01 48.02 39.77 46.98 55.22 48.50
meta-llama/Llama-3.3-70B-Instruct 88.80 77.80 88.40 78.00 83.25 54.37 45.45 60.70 47.39 54.70
meta-llama/Llama-3.1-8B-Instruct 72.20 70.60 77.00 72.40 73.05 51.19 35.23 45.12 33.91 43.20
meta-llama/Llama-3.2-3B-Instruct 65.80 63.60 56.80 57.60 60.95 33.33 28.41 38.84 27.39 33.90
meta-llama/Llama-3.2-1B-Instruct 44.40 41.00 37.60 53.60 44.15 22.22 13.64 29.07 27.39 25.60
Qwen/Qwen3-235B-A22B 96.40 92.40 97.00 89.40 93.80 70.24 72.73 70.93 69.57 70.60
Qwen/Qwen3-30B-A3B 96.60 91.80 97.40 90.20 94.00 64.68 70.45 69.53 56.52 65.40
Qwen/Qwen2.5-72B-Instruct 95.40 89.80 95.60 88.60 92.35 70.63 60.23 61.40 56.09 62.40
Qwen/Qwen3-32B 97.60 94.00 99.00 92.60 95.80 69.05 81.82 68.14 77.83 71.80
Qwen/Qwen3-8B 96.40 93.00 96.20 90.40 94.00 68.65 78.41 73.02 66.52 70.90
Qwen/Qwen3-4B 95.20 91.60 93.60 87.60 92.00 71.03 62.50 75.58 71.74 72.40
Qwen/Qwen3-1.7B 83.20 81.00 80.60 79.60 81.10 48.81 38.64 60.93 41.74 51.50
microsoft/phi-4 92.60 86.40 93.00 85.40 89.35 59.52 57.95 54.19 57.39 56.60
01-ai/Yi-1.5-9B-Chat-16K 90.40 87.40 88.00 85.00 87.70 65.48 63.64 62.09 54.78 61.40
google/gemma-3-1b-it 55.40 56.20 43.00 56.00 52.65 32.14 19.32 33.72 40.87 33.70
3.2 Construction of VerifyBench-Hard
Completion Generation
To construct the dataset, we first generated single-turn completions for the queries described in Section 3.1 using a collection of 18 open-source models. Due to the substantial volume of generations and the associated computational costs, closed-source models were excluded from this stage. In total, we produced approximately 1.45 million completions.
Difficulty Filtering
Next, we employed five top-performing large models on VerifyBench (Llama-3.3-70B-Instruct (Grattafiori et al., 2024), Llama-4-Scout-17B-16E-Instruct (Meta AI, 2025), Qwen2.5-72B-Instruct (Qwen et al., 2025), Qwen3-30B-A3B, and Qwen3-32B (Yang et al., 2025)), which span a diverse range of model sizes and architectures, to evaluate the correctness of the generated completions. Based on their judgments, we identified question-answer-completion tuples exhibiting model disagreement, specifically those for which two models’ assessments diverged from the other three. To ensure balanced and comprehensive representation, we applied stratified sampling across data domains and sources, ultimately selecting 2,000 examples for human annotation.
Human Annotation
We subsequently subjected the selected samples to human annotation, focusing on two key aspects: identifying the answer type and determining the correctness of each completion. Each instance was annotated independently by at least two annotators. In cases where both annotators agreed, the annotation was finalized; when disagreement occurred, a third annotator was consulted to resolve the conflict.
Benchmark Construction
4 Evaluation Results
@@ 4.1 Overall Performance
We evaluate the performance of various verification approaches on both VerifyBench and VerifyBench-Hard. For rule-based baselines, we adopt the widely used math-verify (Kydlíček, 2025) method. In the LLM-as-a-judge setting, we prompt LLMs to perform verification; detailed prompt templates are provided in Appendix C.2. Our evaluation yields several key findings and insights:
Existing models perform well on VerifyBench:
@@
VerifyBench-Hard is challenging:
@@
Small-scale models still have potential for development:
In practical reinforcement learning scenarios, the inference efficiency of the reward system significantly impacts the overall training speed. Since such verification tasks typically involve generative inference, their computational cost is comparable to that of the rollout process itself. Thus, efficiently leveraging smaller models to perform verification is a practical concern worth exploring. According to our results, models with smaller parameters (<3B parameters) exhibit notably poorer performance on VerifyBench, achieving 81.10% accuracy with Qwen3-1.7B and only 60.95% accuracy with Llama-3.2-3B-Instruct, while larger-scale models can achieve over 90% accuracy. Therefore, enhancing the capability of smaller models on these verification tasks represents a valuable direction for future research.

### v3

@@ Abstract
[Uncaptioned image] Project Page https://zju-real.github.io/VerifyBench/
[Uncaptioned image] Benchmark https://huggingface.co/datasets/ZJU-REAL/VerifyBench
[Uncaptioned image] Code https://github.com/ZJU-REAL/VerifyBench
@@ 1 Introduction
In recent years, large language models (LLMs) have exhibited remarkable capabilities, significantly assisting humans across diverse practical domains (DeepSeek-AI et al., 2025b; Grattafiori et al., 2024; Yang et al., 2025). Reinforcement learning from human feedback (RLHF) has been crucial to this progress, with reward models playing a central role by evaluating and scoring model-generated responses to guide training. This approach has led to the development of numerous benchmarks  (Lambert et al., 2025; Liu et al., 2024; Zhou et al., 2024) for systematic reward model evaluation, focusing primarily on pairwise preference judgments between competing responses.
Figure 1: The core distinction between VerifyBench and existing reward benchmarks (Lambert et al., 2025; Liu et al., 2024) is illustrated as follows. Upper panel: Existing reward benchmarks assess the accuracy of a reward system by comparing the ranking of two completions for the same question. Lower panel: In contrast, our proposed VerifyBench evaluates the accuracy of a reward system by determining the correctness of a single completion using a reference answer.
The emergence of specialized large reasoning models (LRMs) (DeepSeek-AI et al., 2025a; Qwen Team, 2024; Kimi Team et al., 2025) such as OpenAI’s o1(OpenAI, 2024) and DeepSeek-R1(DeepSeek-AI et al., 2025a) has fundamentally changed this landscape. These models achieve unprecedented performance on reasoning tasks through specialized reinforcement learning techniques that differ from standard RLHF approaches. A key distinction in training methodologies for LRMs is their reliance on reference-based reward systems, where rewards are assigned based on alignment between model-generated responses and authoritative reference answers. This approach has been implemented variously across leading models, with Deepseek-R1 employs a rule-based reward to prevent reward hacking, whereas models like Seed1.5-Thinking (Seed et al., 2025) adopt model-based reward systems to generate more precise and robust signals.
Despite the widespread adoption of reference-based reward systems in training state-of-the-art reasoning models, a significant gap exists in our ability to evaluate these systems systematically. Current benchmarks focus almost exclusively on preference-based evaluation, assessing rewards on their ability to rank competing responses correctly. This approach fails to capture the requirements of reference-based verification, where responses must be judged against objective ground truths rather than relative preferences (Kim et al., 2024). The absence of dedicated benchmarks for reference-based reward systems has limited researchers’ ability to assess, compare, and improve their verification methodologies effectively, potentially impeding progress in reasoning model development.
@@ 2 Preliminaries
$\text{Accuracy}=\frac{1}{|D|}\sum_{(q,r_{w},r_{l})\in D}\mathbb{I}\left[R_{\varphi}(q,r_{w})>R_{\varphi}(q,r_{l})\right]$ (2)
@@
where $q$ denotes the user-issued query, $gt$ denotes the ground-truth reference answer, $r$ represents the model-generated response, and $\varphi$ encapsulates either the learned parameters of the reward model or the established evaluation criteria used to assess the alignment among $q$, gtgt, and $r$.
@@
$\text{Accuracy}=\frac{1}{|D|}\sum_{(q,gt,r,y)\in D}\mathbb{I}\left[\mathbb{E}(R_{\varphi}(q,gt,r))=y\right]$ (4)
@@ 3 Benchmark Construction
In this paper, we introduce two benchmarks, VerifyBench and VerifyBench-Hard, to evaluate reference-based reward systems. The VerifyBench benchmark (Section 3.1) is designed to reflect naturally distributed data, whereas VerifyBench-Hard (Section 3.2) comprises samples exhibiting high levels of disagreement among models, thereby assessing a model’s ability to provide reliable judgments in ambiguous or challenging scenarios.
@@ 3.1 Construction of VerifyBench
To emulate realistic reinforcement learning (RL) scenarios involving reference-based reward systems, we curate a comprehensive collection of open-source reasoning problems paired with corresponding reference answers. These problems encompass three primary categories, general reasoning, logical reasoning and mathematical reasoning, and are aggregated from 41 distinct sources. A complete list of these data sources is provided in Appendix C.
@@
To comprehensively evaluate model performance across diverse answer formats, we define four canonical answer types: numerical values, algebraic expressions, multiple-choice selections, and free-form strings. Utilizing a general-purpose LLM Llama-3.3-70B-Instruct (Grattafiori et al., 2024), we performed automatic answer-type classification with a prompt (Appendix  E.1). Questions that fall outside these categories, such as proof-based or open-ended prompts, were excluded from further analysis. Following classification, we randomly sampled 2,000 instances per answer type, resulting in a final candidate pool of 8,000 questions.
@@
We subsequently employed 22 widely used open-source and proprietary models (see Appendix D) to generate single-shot completions for the curated set of 8,000 questions, resulting in a total of 176,000 completions. To assign initial correctness labels, we utilized Llama-3.3-70B-Instruct (Grattafiori et al., 2024) within a prompt-based judgment framework. For each question, we randomly selected four completions, two labeled as correct and two labeled as incorrect by the model, and retained them for subsequent human annotation.
@@
We conducted human annotation for the aforementioned questions and their associated completions. The annotation procedure comprised two primary tasks: (1) identifying the most appropriate answer type for each question based on its formulation and corresponding ground-truth answer, and (2) evaluating the correctness of each of the four completions. Each question was independently annotated by at least two annotators. If their annotations were consistent, the labeling was finalized; otherwise, a meta-annotator resolved disagreements to ensure consistency and finalize the labels. A detailed description of the human annotation process is provided in Appendix F.
@@
Following human annotation, we identified notable biases in the models’ predictions regarding both answer types and completion correctness, leading to imbalanced data distributions. To mitigate this issue, we performed controlled downsampling to ensure uniform category-level representation and balanced correctness labels. Specifically, we retained 250 questions per answer type, resulting in a total of 1,000 questions. Each question is paired with exactly two completions, one correct and one incorrect. The resulting dataset, VerifyBench, thus comprises 2,000 well-balanced question-answer-completion-correctness tuples. Detailed statistics for VerifyBench are provided in Table 1, and illustrative examples are presented in Appendix I.
Table 1: Benchmark statistics of VerifyBench (VB) and VerifyBench-Hard (VB-H).
@@
Answer Type Distribution
@@
3.2 Construction of VerifyBench-Hard
Completion Generation
To construct the dataset, we first generated single-turn completions for the queries described in Section 3.1 using a collection of 18 open-source models (Appendix D). Due to the substantial volume of generations and the associated computational costs, closed-source models were excluded from this stage. In total, we produced approximately 1.45 million completions.
Difficulty Filtering
Next, we employed five top-performing LLMs on VerifyBench (Llama-3.3-70B-Instruct (Grattafiori et al., 2024), Llama-4-Scout-17B-16E-Instruct (Meta AI, 2025), Qwen2.5-72B-Instruct (Qwen et al., 2025), Qwen3-30B-A3B, and Qwen3-32B (Yang et al., 2025)), which span a diverse range of model sizes and architectures, to evaluate the correctness of the generated completions. Based on their judgments, we identified question-answer-completion tuples exhibiting model disagreement, specifically those for which two models’ assessments diverged from the other three. To ensure balanced and comprehensive representation, we applied stratified sampling across data domains and sources, ultimately selecting 2,000 examples for human annotation.
Human Annotation
We subsequently subjected the selected samples to human annotation, focusing on two key aspects: identifying the answer type and determining the correctness of each completion. Each instance was annotated independently by at least two annotators. In cases where both annotators agreed, the annotation was finalized; when disagreement occurred, a meta-annotator was consulted to resolve the conflict. A detailed description of the human annotation process is provided in Appendix F.
Benchmark Construction
4 Evaluation Results
@@
gpt-4o-2024-11-20 94.80 90.20 96.80 90.80 93.15 71.43 65.91 75.35 71.30 72.60
gpt-4o-mini 95.80 89.80 95.80 90.00 92.85 69.05 72.73 74.19 72.17 72.30
DeepSeek-V3-0324 96.80 93.00 97.60 91.60 94.75 81.35 77.27 81.63 75.22 79.70
DeepSeek-R1-0528 98.00 92.60 98.00 92.00 95.15 82.14 81.82 90.93 85.22 86.60
gpt-oss-20b 97.40 94.20 98.40 89.20 94.80 82.54 79.55 90.23 76.96 84.30
gpt-oss-120b 98.00 94.80 99.20 91.40 95.85 84.13 80.68 92.56 86.09 87.90
Llama-4-17B-16E-Instruct 94.20 86.80 89.80 89.25 90.01 48.02 39.77 46.98 55.22 48.50
Llama-3.3-70B-Instruct 88.80 77.80 88.40 78.00 83.25 54.37 45.45 60.70 47.39 54.70
Llama-3.1-8B-Instruct 72.20 70.60 77.00 72.40 73.05 51.19 35.23 45.12 33.91 43.20
Llama-3.2-3B-Instruct 65.80 63.60 56.80 57.60 60.95 33.33 28.41 38.84 27.39 33.90
Llama-3.2-1B-Instruct 44.40 41.00 37.60 53.60 44.15 22.22 13.64 29.07 27.39 25.60
Qwen3-235B-A22B 96.40 92.40 97.00 89.40 93.80 70.24 72.73 70.93 69.57 70.60
Qwen3-30B-A3B 96.60 91.80 97.40 90.20 94.00 64.68 70.45 69.53 56.52 65.40
Qwen2.5-72B-Instruct 95.40 89.80 95.60 88.60 92.35 70.63 60.23 61.40 56.09 62.40
Qwen3-32B 97.60 94.00 99.00 92.60 95.80 69.05 81.82 68.14 77.83 71.80
Qwen3-8B 96.40 93.00 96.20 90.40 94.00 68.65 78.41 73.02 66.52 70.90
Qwen3-4B 95.20 91.60 93.60 87.60 92.00 71.03 62.50 75.58 71.74 72.40
Qwen3-1.7B 83.20 81.00 80.60 79.60 81.10 48.81 38.64 60.93 41.74 51.50
phi-4 92.60 86.40 93.00 85.40 89.35 59.52 57.95 54.19 57.39 56.60
Yi-1.5-9B-Chat-16K 90.40 87.40 88.00 85.00 87.70 65.48 63.64 62.09 54.78 61.40
gemma-3-1b-it 55.40 56.20 43.00 56.00 52.65 32.14 19.32 33.72 40.87 33.70
Model-based Verifier
xVerify-0.5B-I 79.60 74.40 88.80 87.20 82.50 60.32 62.50 91.86 77.83 78.10
xVerify-3B-Ia 89.80 86.40 95.00 90.20 90.35 73.81 78.41 93.49 73.91 82.70
xVerify-8B-I 93.80 89.60 95.20 89.20 91.95 69.05 76.14 93.49 81.74 83.10
Tencent-Qwen2.5-7B-RLVR 96.80 95.40 93.00 90.80 94.00 75.43 75.00 61.36 63.64 72.20
CompassVerifier-3B 92.00 84.80 94.60 88.60 90.00 77.38 84.09 93.95 75.22 84.60
CompassVerifier-7B 92.80 88.60 97.20 88.40 91.75 78.97 85.23 95.35 76.09 85.90
CompassVerifier-32B 94.20 89.80 96.80 89.40 92.55 80.56 82.95 95.58 83.48 87.90
@@ 4.1 Overall Performance
We evaluate the performance of various verification approaches on both VerifyBench and VerifyBench-Hard. For rule-based baselines, we adopt the widely used math-verify (Kydlíček, 2025) method. In the LLM-as-a-judge setting, we prompt LLMs to perform verification; detailed prompt templates are provided in Appendix E.2. Additionally, we conduct evaluations of existing model-based verifiers, employing prompts consistent with their provided templates to ensure fairness (Chen et al., 2025; Su et al., 2025; Liu et al., 2025a). Our evaluation yields several key findings:
Existing models perform well on VerifyBench.
@@
VerifyBench-Hard is challenging.
@@
Small-scale models still have room for development.
In practical reinforcement learning scenarios, the inference efficiency of the reward system significantly impacts the overall training speed. Since such verification tasks typically involve generative inference, their computational cost is comparable to that of the rollout process itself. Thus, efficiently leveraging smaller models to perform verification is a practical concern worth exploring. According to our results, models with smaller parameters (<<3B parameters) exhibit notably poorer performance on VerifyBench, achieving 81.10% accuracy with Qwen3-1.7B and only 60.95% accuracy with Llama-3.2-3B-Instruct, while larger-scale models can achieve over 90% accuracy. Therefore, enhancing the capability of smaller models on these verification tasks represents a valuable direction for future research.
@@
4.2 Reference-answers Play an Important Role in Verification
The benchmark proposed in this work fundamentally differs from existing reward benchmarks by explicitly incorporating reference answers, thereby aligning more closely with the training setups of contemporary reasoning LLMs. To isolate the impact of reference answers on verification performance, we conduct an ablation study in which models are evaluated without reference answers provided in the prompt; the prompt format used is detailed in Appendix E.3.
Experimental results, summarized in Table 3, reveal a performance degradation of approximately 5-18% when reference answers are excluded. These findings underscore the crucial role of reference answers in reasoning-oriented RL, suggesting they provide a more reliable and informative supervision signal during reward modeling.
4.3 Performance of Reference-free Reward Models
Our experimental results show that VerifyBench introduces a level of challenge comparable to existing reward benchmarks, with the absence of reference answers. Reference-free reward models achieve sub-80% accuracy on VerifyBench, highlighting its difficulty. Furthermore, domain-specific reward models exhibit inferior performance on general reward benchmarks (Liu et al., 2024; Lambert et al., 2025) compared to VerifyBench, validating the benchmark’s design objectives.